### v1

@@ 6.2.2 Learning methods
Distance metric learning Distance Metric Learning (DML) aims to learn a discriminative embedding in which similar samples are closer together, and dissimilar samples are separated [52]. DML emerged from the concept of contrastive loss, which turns this principle into a learning objective [16]. Unlike classification losses, which learn a class logit vector, the contrastive loss in DML captures the relationships among samples: it trains a Siamese Network, which consists of two identical subnetworks whose architecture, configurations, and weights are the same, to predict whether two inputs are from the same class. This is achieved by putting their embedding close to each other (for the same class) and far apart (for different classes) [16, 52].
In particular, DML based approach can be effective to overcome the issue of poor generalization performances of DL models caused by annotation bias problems. For instance, Zhang et al. [229] used similar finding retrieval based on DML to improve DL models generalizability. Their method requires an extra "clean" dataset with pathological-proven labels to re-label a noisy dataset according to similar finding retrieval. First, a Siamese Network is trained on randomly selected pairs of images from the clean dataset. Each pair is assigned a label depending on whether the two images of this pair belong to the same class. The model is trained under the assumption that if two images belong to the same class, their feature vectors should be close in the feature space, and if they belong to different classes, their feature vectors should be far apart. To enforce this, a contrastive loss is applied over the Euclidean distance metric (similarity score) induced by the malignancy representation. Next, during the re-labeling phase, each sample from the clean dataset used in training is paired with an "under-labeled" sample from the noisy dataset. These noisy samples are then sorted based on their similarity scores with their corresponding clean sample. The new label for each sample from the noisy dataset is determined by averaging the labels of the top 20% of its partner samples in the ranking list of similarity scores. This process aims to provide more accurate labels for the noisy label samples based on their similarity to pathologically confirmed samples in the clean dataset. In addition, taking the average labels of top-ranked similarity samples is intended to increase the confidence of label propagation.
Zhang et al. [229] applied the aforementioned method using the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [9], a prominent source of public dataset of thoracic Computed Tomography (CT) image data. Despite its popularity, the LIDC-IDRI dataset is considered a noisy dataset because of the absence of clinical information. More precisely, four qualified thoracic radiologists assessed nodule characteristics and provided a rating of malignancy rating (from 1 to 5). A thorough examination of the dataset revealed that the ’likelihood of malignancy’ characteristic was subjective since it was established based on the assumption of a 60-year-old male smoker, due to the absence of any clinical information about the patients. To challenge the low confidence labels of LIDC, the authors proposed to use a pathological-proven labeled dataset, the SCH-LND [230] dataset. This extra dataset includes 180 solid nodules (90 benign / 90 malignant). The idea is to apply this dataset to transfer pathological knowledge for various clinical indications. To this end, two re-labeling methods were explored to obtain new ground-truth labels on the LIDC database. The first method produces the malignancy label using a machine annotator: a nodule classifier that was pre-trained on LIDC data and fine-tuned on SCH-LND data. The second method uses a machine comparator to rank the top nodules’ labels: a metric-based network computes the correlation between pairs of nodules. According to this study, relabeling through metric learning outperformed the general supervised model. It suggests that the input pairs produced by random sampling provide a data augmentation effect to learning with limited data.
@@
With the aim of being robust to inherent observer-variability in training labels, Van Woudenberg et al. [195] also employed the DML based approach in another form within the differential learning approach. The differential learning consists of training the model on an auxiliary comparison task, i.e., determining whether a clinical parameter is significantly different between two patients. This task is considered to be an easier task and less prone to subjectivity. It is modeled using a Siamese Network to compare the estimated clinical parameter that are sufficiently apart from each other based on the generated representations. The proposed approach showed good results in the assessment of left ventricle measurements in echocardiography cine series. In this context, differential learning was integrated as an auxiliary task by computing whether there is a significant difference in Ejection Fraction (EF) between two patients (normal vs severe EF). It showed enhanced performances when evaluated on 1) a large cart-based dataset consisting of 28,577 echo cines obtained from 23,755 patients and 2) 51 echo cines acquired from 23 heart-failure patients using a Point-Of-Care Ultrasound (POCUS).
@@
While the aforementioned methods tackled the problem of noise in labels in form of subjective interpretations between readers, Seibold et al. [168] addressed inconsistent labels generated from unstructured medical reports via text classifiers, typically by using classifiers which may not perfectly capture the nuances of the medical terminology and context. The authors argued that traditional methods rely on a predefined set of categories extracted from medical reports which can cause some limitations and inaccuracies. To overcome this problem, they proposed a contrastive language-image pre-training on report-level approach to learn directly from the unstructured medical reports. They applied a contrastive global-local dual-encoder architecture to learn concepts directly from unstructured medical reports while maintaining its ability to perform free-form classification. In contrast to the previously mentioned approach, DML is combined with self-supervision, by integrating SimSiam [35]. In this framework, two augmented versions of the same input image are created. These images are then processed through a backbone network, a three-layer encoder-head, and a two-layer prediction-head, to enforce similarity between the two views. Furthermore, they use the augmented images from the pre-training objective to mirror their text-image objectives to the augmented samples. Their approach performed on par with direct label supervision approach when evaluated on the large-scale chest X-ray datasets MIMIC-CXR [90], CheXpert [85], and ChestX-Ray14 [207] for disease classification.
@@
Finally, similar to [168], Kurian et al. [102] employed a DML method in conjunction with a self-supervised approach. Their method is based on a contrastive learning framework and feature aggregating memory banks to identify and increase the emphasis on clean training samples. In essence, it consists of 1) a warm-up phase that includes contrastive learning following a weight update based on cross-entropy, 2) a weight calculation phase in which the weights of the training samples are adjusted in accordance with their similarity to the features preserved in the memory bank and 3) a final classification training phase which consists of training the model using a weighted cross entropy loss function. The method uses ResNet-18 neural architecture trained using SimCLR, a self-supervised training. In the warm-up phase, training involves both cross-entropy loss and contrastive loss. A maximum loss miner function, using the contrastive loss, identifies ‘hard-pairs’ and noisy labels based on cosine similarity. The subset with maximum contrastive loss represents the hard pairs, all the feature vectors except the hard-pairs are updated class-wise into a memory bank, which is limited to a fixed size. The features are thus appended in this memory bank in a FIFO manner. In the weight calculation phase, sample weights are computed using cosine similarities with memory bank features of the same class. Next, a K-medoids computation for the features in the memory bank is carried out to evaluate the cosine similarity between the samples and the medoids. Exponential of the similarity scores are normalized between 0 and 1. In the final classification training phase, these scores are applied as weights for the cross-entropy loss. The authors demonstrated the effectiveness of their approach when evaluated on two histology datasets: 1) the ICIAR 2018 Grand Challenge dataset called Breast Cancer Histology (BACH) dataset [7] and 2) the basal versus Luminal A PAM50 subtype classification from H&E stained WSI extracted from TCGA-BRCA dataset.
@@
Active learning Active learning can be combined with noisy labels for improving DL performances.
In this context, Son et al. [182] proposed this strategy [^25] https://bitbucket.org/woalsdnd/codes_and_data to enhance the model’s learning efficiency for detecting a rare disease. Their method involved assigning "normal" pseudo-labels to a large number of publicly available unlabeled images. This set was combined with a small set of labeled images (having the targeted rare disease) for initial training. The authors thus introduced noise in the pseudo-labels as it is likely that some of the images pseudo-labeled as "normal" actually contain the targeted disease. After the CNN is initially trained on the pseudo-labeled dataset, it is used to identify images that it predicts as rare disease with high confidence (greater than 0.5). This step effectively filters out a significant portion of the noise by focusing on cases where the model has high confidence, thereby reducing the need to manually inspect a vast number of images. This process significantly narrows down the number of images to be manually reviewed for the rare disease. The active learning process allows to screen for the positive selected cases. It helps in correcting the initial noise introduced by pseudo-labeling. The refined dataset is then used for training. This strategy attained an AUC of 0.9993 on the PALM competition, ranking first on the off-site validation set.
@@
Zero-shot learning Zero-shot learning (ZSL) is a technique enabling machine learning algorithms to recognize objects belonging to new, previously unseen classes, using semantic descriptions. ZSL fundamentally relies on auxiliary information about seen and unseen classes in the form of semantic description. A key advancement in this field is the Generalized Zero-Shot Learning (GZSL), where the test data may originate from either seen or unseen classes, posing a more realistic and challenging scenario. In ZSL, the process involves mapping a visual feature vector extracted from an input image to a semantic space created using auxiliary information. This mapping is crucial for bridging the gap between the visual and semantic domains, enabling the recognition of unseen classes.
Based on these principles, Paul et al. [143] proposed a GZSL for the diagnosis of chest radiographs. Their approach is based on a Multi-View Semantic Embedding (MVSE) network, which integrates three distinct semantic spaces: one derived from X-ray reports, another from CT radiology reports, and a third based on visual traits used by radiologists. The inclusion of the CT semantic space is particularly insightful, as most diseases identifiable from chest X-rays are also diagnosable from chest CTs, which provide richer information. The methodology involves a two-branch autoencoder for performing semantic embeddings into these X-ray and CT semantic spaces. Each branch is supplemented with a novel guiding network that leverages the trait-based semantic space, thereby enhancing the model’s ability to interpret complex medical images. To improve performance for unseen classes, the framework incorporates a self-training strategy. This process begins by creating a self-training set comprising unlabeled X-ray images from both seen and unseen classes. The self-training is executed in two steps: initial inference and model fine-tuning. Initially, class probabilities for unlabeled images from the self-training set are computed using the trained MVSE network. Images are then selected for both seen and unseen classes based on the highest confidence scores. Subsequently, the model is fine-tuned with this selectively chosen data for each class. This refined model is then deployed for generalized zero-shot diagnosis of chest X-rays. During testing, for a given X-ray image that may belong to either seen or unseen classes, the model computes distances in both the X-ray and CT semantic spaces from the respective class signatures. The final class probability is determined by dynamically balancing the importance assigned to the X-ray and CT branches. Then authors showed that their model achieved robust generalization capabilities when tested on NIH Chest X-ray dataset (NIH), hand-labeled subset of NIH dataset (NIH-900), Open-i dataset, PubMed Central dataset (PMC) and the CheXpert dataset.
@@
The Balanced Gradient Contribution (BGC) strategy is a training approach designed to manage the significant statistical differences between domains [158]. This method addresses the issue of large variance in gradients that occurs due to the distinct nature of data from each domain. In the context of domain generalization, the BGC method could be employed to balance the learning from different domains by adjusting the contribution of gradients from each domain during the training process.
This strategy was integrated in the work of Elbatel et al. [53]. They proposed a Seamless Iterative Semi-supervised correction of imperfect labels (SISSI) [^26] https://github.com/marwankefah/SISSI, a novel framework for training object detection models with noisy and missing annotations. They introduced a range of image processing and deep learning methods to make iterative label correction. They applied a domain adaptation strategy to address noisy labels by exploiting a source labelled dataset (Cellpose), to enhance training on a target noisy dataset (CellLab). In the early phase, they used a mixed-batch training combining both training datasets to train a Faster R-CNN mode using BGC [158], which generates stable gradient directions. To determine the end of this phase, they used ADELE method which permits to detect when the network starts memorizing the initial noisy annotations. Next, they applied a semi-supervised phase for each cycle. They used label correction strategy. To this end, they used test-time augmentation and weighted box fusion techniques to filter potential bounding boxes and hence producing confident bounding boxes. They also proposed to use synthetic like image adaptation according to pseudo labels. This method would solve the problem of undetected cells in the pseudo labels which would affect the training negatively. Thus, the authors generated dynamically synthetic-like images for continual training.
@@
Robust loss function Robust loss functions focus mainly on improving the loss to build robust DL network.
To address the problem of model overfitting due to label ambiguity and noisy labels, Sun et al. [187] proposed to use deep log-normal label distribution learning and focal loss. The concept of deep log-normal label distribution learning is inspired by the paradigm of label distribution learning [61, 58], where an instance is assigned a label distribution, aiming to learn a mapping from instance to label distribution. For the targeted application of pneumoconiosis staging on chest radiographs, the construction of the label distribution is to estimate the tolerance of each stage of pneumoconiosis. The authors modeled the label distribution using asymmetric log-normal distribution.
@@
To enforce label distribution learning, the KL Divergence loss is employed using the label distribution and the prediction of the staging network. This loss aims to measure the distance between two continuous distributions: the label distribution ($y^{d}$) and the network prediction after the Softmax function.
In addition, the authors added a regularization term, the cross-entropy loss, to strengthen the learning abilities of the model on unambiguous samples and deal with subjective inconsistency. Combining both loss, the KL divergence loss and the cross-entropy loss, constitute the focal staging loss. However, using the aforementioned loss together results in an optimization inconsistency. To solve this issue, an instance-level drop parameter was introduced to skip samples with better predicted results during the optimization process.
Zhu et al. [241] employed the focal loss to enhance the training process by emphasizing hard samples. It is defined as follows:
@@
where $q(y|x)$ is the predicted probability and $\gamma$ is a hyperparameter. To reduce the relative loss for well-classified examples and put more focus on hard, misclassified examples, the parameter $\gamma$ in the focal loss function was set to 2. This technique aims at improving the robustness of the model against noisy labels and ensuring that the model learns effectively from the most challenging parts of the data.
Hu et al. [81] proposed a robust training method for solving the classification problem with imbalanced data. The targeted application was the diagnosis from imbalanced CT images. Their framework, called Deep Supervised Network with a Self-Adaptive Auxiliary Loss (DSN-SAAL), consisted of integrating a novel loss function. The loss function tackles both the effects of data overlap between CT slices and noisy labels. In order to consider the effects of data overlap between CT slices, they propose adjusting the weight of samples in the Cross-Entropy (CE) loss function as follows:
@@
where $q(y|x)$ denotes the output of the classifier and $p(y|x)$ represents the ground-truth distribution over labels for sample $x$. $k_{i}$ is the number of samples in the $k^{th}$ class and $\alpha$ is a learnable parameter which represents the effective sample factor to measure the ratio of the effective number of samples. In addition, they introduced a regularization term, called Reverse Cross Entropy (RCE) to combat noisy labels:
@@
In this setting, since some labels are incorrect and $p(y|x)$ does not represent the true class distribution, they employed $q(y|x)$ as the ground truth and $p(y|x)$ as the class probability of the outputs, hence the name reverse cross-entropy. Finally, the proposed self-adaptive auxiliary loss can be formulated by combining the aforementioned losses (Equations 26 and 27) while adding a weighting hyperparameter $\beta$ to balance between them. The authors showed that their approach outperformed the state-of-the-art methods when evaluated on COVID19-Diag and three public COVID-19 diagnosis datasets.
To handle the subjectivity of annotation, Yu et al. [226] proposed a Grading Cross Entropy (GCE) loss. The loss is designed to account for the feature continuity of disease grades and the inherent progression of disease, recognizing that misclassifications are more likely to occur between adjacent grades rather than distant ones. It is defined as follows:
@@
where $p(y_{i}|x)$ is the $i$-th element of the one-hot encoded label of the input $x$, $\mathbb{N}(i)$ the neighboring indexes of grade $i$, $q(y_{j}|x)$ denotes the $j$-th element of the model predictions and $w_{ij}$ represents the weight of grade $j$ with the annotated label $i$. This weighting system allows the GCE loss to be more flexible than the CE loss in handling noise by setting different weights to neighboring labels and the annotated label. Therefore, GCE can effectively reduce overfitting to noisy data.
Hermoza et al. [78] tackled the problem of predicting survival time from medical images using both censored and uncensored data. They proposed an Early-Learning Regularization (ELR) loss regularization loss to handle pseudo label noise. ELR is designed to mitigate the impact of noisy pseudo labels in the survival time prediction setting. It is used to ensure that training continues for samples where the model’s prediction aligns with the temporal ensembling momentum (considered "clean" pseudo-labeled samples) and is minimized for noisy pseudo-labeled samples. The temporal ensembling momentum is a blend of the current and previous predictions. The ELR loss is defined as follows:
@@
where $\tilde{z}_{i}^{c(e)}=\phi\tilde{z}_{i}^{c(e-1)}+(1-\phi)z_{i}^{c(e)}$ is the temporal ensembling momentum of the prediction with $e$ denoting the training epoch and $\phi\in[0,1]$. $\sigma$ is the sigmoid function, $z^{c}$ is the output of the model, and $c$ denotes the number of classes.
Liu et al. [121] proposed a similar approach based on a new training module called Non-Volatile Unbiased Memory (NVUM) [^27] https://github.com/FBLADL/NVUM, which non-volatility stores running average of model logits. The authors proposed a regularization loss based on this memory module to minimize the differences between the model’s current logits and those from its initial learning phase. In essence, the model $f_{\theta}$ is trained on noisy labeled dataset using binary cross-entropy loss. In addition to the binary cross-entropy loss, they integrated the following regularization term:
@@
where $\tilde{z}^{c}$ is a memory module that stores an unbiased multi-label running average of the predicted logits of all training samples and $\tilde{z}^{c}$ employs the class prior distribution $\pi$ for updating, denoted by $\pi(j)=\frac{1}{n}\sum_{i=1}^{n}\hat{y}(j)$ for $j\in\{1,...,c\}$, where $n$ is the total number of sample in the noisy dataset, and $\hat{y}$ denotes a one-hot noisy label.
Initially, $\tilde{z}$ is initialized with zeros, it is updated in every epoch using the following formula:
@@
where $\beta\in[0,1]$ is a hyperparameter controlling the volatility of the memory storage. $\beta$ is set to $0.9$ representing a non-volatile memory. Since in the early stage of training the model is considered to be robust to noisy label, this regularization enforces consistency between the current model logits and the logits produced at the beginning of the training. The NVUM approach was evaluated on new benchmarks, where training is performed on noisy multi-label imbalanced chest X-ray (CXR) training sets, formed by Chest-Xray14 and CheXpert, and the testing is performed on the clean multi-label CXR datasets OpenI and PadChest. Their approach outperformed previous state-of-the-art CXR classifiers and previous methods that can deal with noisy labels on all evaluations.
@@
Shi et al. [175] proposed a semi-supervised deep learning approach that leverages the semantic information of annotated images and explores the information hidden in unlabeled data, while being robust to noisy labels. Their approach is based on self-ensembling based deep architecture. It is inspired by Temporal Ensembling (TE) [105], which produces an ensemble target for each label prediction of training samples. This is done by applying Exponential Moving Average (EMA) to the predictions within several previous training epochs and then minimizing the difference between the label prediction and its ensemble target. Shi et al. [175] proposed Graph Temporal Ensembling (GTE) to improve TE. GTE creates ensemble targets for feature and label predictions of each training sample and forms consensus predictions under different configurations. Using a graph, the ensemble targets within the same class are mapped into a cluster so that they are further enhanced. Next, a consistency loss is utilized to form consensus predictions under different configurations.
$\begin{split}\mathcal{L}=&\frac{\tau(t)}{c|B|}\sum_{i,j\in B}\bigg{(}\lambda_{1}\left\|z_{i}^{f}-\sum_{j=1}^{|B|}w_{ij}\hat{z}_{j}^{f}\right\|^{2}_{F}\\ &+\lambda_{2}\left\|z_{i}^{c}-\sum_{j=1}^{|B|}w_{ij}\hat{z}_{j}^{c}\right\|^{2}_{F}\bigg{)},\end{split}$ (32)
where $B$ denotes the batch of selected images, $|B|$ represents the number of selected images, $c$ is the number of classes. The prediction for image $i$ is given by $z_{i}^{c}$, while $y_{i}$ is the true label. The time-dependent weighting function to gradually enhance the weight of the consistency cost is represented by $\tau(t)$, and $\lambda_{1}$ and $\lambda_{2}$ are non-negative constants that balance the terms in the loss function. The feature representation for image $i$ is $z_{i}^{f}$, and $\tilde{z}_{j}^{f}$ is obtained by applying EMA to feature representations within multiple previous training epochs, i.e., $\hat{z}_{j}^{f}=\frac{\tilde{z}_{j}^{f}}{(1-\alpha^{t})}$, where $\tilde{z}_{i}^{f}=\alpha\tilde{z}_{i}^{f}+(1-\alpha)z_{i}^{f}$ where $\alpha$ is an ensembling momentum term to control how far the ensemble $\tilde{z}_{j}^{f}$ reaches into the training history, and $t$ is the current number of training epochs. Similarly, $\tilde{z}_{j}^{c}$ is obtained by applying EMA to label predictions within multiple previous epochs, i.e., $\hat{z}_{j}^{c}=\frac{\tilde{z}_{j}^{c}}{(1-\alpha^{t})}$, where $\tilde{z}_{i}^{c}=\alpha\tilde{z}_{i}^{c}+(1-\alpha)z_{i}^{c}$. The weighting factor between images $i$ and $j$ is $w_{ij}$. The authors validated the proposed method with extensive experiments on lung and breast cancer datasets that contain thousands of images.

### v3

@@ 5.2.2 Learning methods
Distance metric learning Distance Metric Learning (DML) aims to learn a discriminative embedding in which similar samples are closer together, and dissimilar samples are separated [147]. DML emerged from the concept of contrastive loss, which turns this principle into a learning objective [148]. The contrastive loss in DML captures the relationships among samples: it trains a Siamese Network, which consists of two identical subnetworks whose architecture, configurations, and weights are the same, to predict whether two inputs are from the same class. This is achieved by putting their embedding close to each other (for the same class) and far apart (for different classes) [148, 147].
Zhang et al. [85] used similar finding retrieval based on DML to improve DL models’ generalizability. They employed an extra “clean” dataset with pathological-proven labels (the SCH-LND [149] dataset) to re-label a noisy dataset, the Lung Image Database Consortium and Image Database Resource Initiative (LIDC-IDRI) [150]. Two re-labeling methods were explored: 1) a nodule classifier pre-trained on LIDC data and fine-tuned on SCH-LND data for malignancy labeling, and 2) a metric-based network (Siamese Network) to rank top nodule labels by computing correlations between nodule pairs. The Siamese Network was trained on randomly selected pairs of images from the clean dataset using the contrastive loss. During the re-labeling phase, each sample from the SCH-LND dataset was paired with an “under-labeled” sample from LIDC, and these under-labeled samples were sorted based on their similarity scores. The new label for each sample in the LIDC dataset was obtained by averaging the labels of the top 20% of its partner samples with the highest similarity scores. According to this study, relabeling through metric learning outperformed the general supervised model, suggesting that the input pairs produced by random sampling provide a data augmentation effect to learning with limited data.
@@
Van Woudenberg et al. [86] employed a DML-based approach within the differential learning approach to address observer-variability in training labels. This method involved training the model on an auxiliary comparison task – determining whether a clinical parameter differed significantly between two patients– considered easier task and less subjective. A Siamese Network was employed to compare the estimated clinical parameter based on the generated representations. The approach showed good results in assessing left ventricle measurements in echocardiography cine series. Differential learning was integrated as an auxiliary task by computing whether there is a significant difference in Ejection Fraction (EF) between two patients (normal vs severe EF). It showed enhanced performances when evaluated on two datasets: a large cart-based dataset consisting of 28,577 echo cines obtained from 23,755 patients and 51 echo cines acquired from 23 heart-failure patients using a Point-Of-Care Ultrasound.
@@
While previous methods addressed noisy labels due to subjective interpretations, Seibold et al. [87] tackled inconsistent labels generated from unstructured medical reports via text classifiers. They proposed a contrastive language-image pre-training on report-level approach using a global-local dual-encoder architecture to learn concepts directly from unstructured medical reports and perform free-form classification. Unlike previously mentioned methods, they combined DML with self-supervision by integrating SimSiam [151]. Two augmented versions of the same input image were created and then processed through a backbone network, an encoder-head, and a prediction-head to enforce similarity between the two views. Furthermore, they use the augmented images from the pre-training objective to mirror their text-image objectives to the augmented samples. This approach matched the performance of direct label supervision on large-scale chest X-ray datasets (MIMIC-CXR [126], CheXpert [128], and ChestX-Ray14 [127]) for disease classification.
@@
Kurian et al. [88] employed a DML method combined with Self-Supervised Learning (SSL), based on a contrastive learning framework and feature aggregating memory banks. The method comprises three phases. The first phase, the warm-up phase, uses both cross-entropy loss and contrastive loss. A maximum loss miner function, using the contrastive loss, identifies ‘hard-pairs’ and noisy labels based on cosine similarity. The subset with maximum contrastive loss represents the hard pairs, while all the other feature vectors are updated class-wise into a fixed-size memory bank. The second phase, the weight calculation phase, assigns weights to each training sample based on their cosine similarity to features in the memory bank, which represent clean samples. K-medoids for the features in the memory bank were also found to compute the cosine similarity for the samples with the medoids. The third phase, the final classification training phase, involves training the model with a weighted cross-entropy loss, applying the computed similarity scores as weights.
@@
Active learning Active learning can be combined with noisy labels to enhance DL performances.
Son et al. [70] proposed a strategy [^26] https://bitbucket.org/woalsdnd/codes_and_data to improve the detection of a rare disease by assigning “normal” pseudo-labels to a large number of publicly available unlabeled images. This set was combined with a small set of labeled images with the targeted rare disease for initial training. Noise was introduced in the pseudo-labels since some of the pseudo-labeled “normal” images likely contained the disease. Initially, their model was trained on the pseudo-labeled dataset. It was then used to identify rare disease images with high confidence predictions (greater than 0.5), effectively filtering out noise by focusing on cases where the model has high confidence. This process significantly reduces the number of images to be manually reviewed for the rare disease. The active learning process allows to screen for the positive selected cases and correct the initial noise introduced by pseudo-labeling. The refined dataset was used for final training, achieving an AUC of 0.9993 on the PALM competition, ranking first on the off-site validation set.
@@
Zero-shot learning Zero-shot learning (ZSL) is a technique enabling machine learning algorithms to recognize objects belonging to new, unseen classes, with the help of semantic descriptions. A pragmatic version of ZSL is the Generalized Zero-Shot Learning (GZSL), where the test data may originate from either seen or unseen classes.
Paul et al. [89] proposed a GZSL for the diagnosis of chest radiographs using a Multi-View Semantic Embedding (MVSE) network, integrating semantic spaces from X-ray reports, radiology reports, and visual traits used by radiologists. They employed a two-branch autoencoder for semantic embeddings into X-ray and CT semantic spaces. Each branch was supplemented with a guiding network leveraging the trait-based semantic space. To improve performance for unseen classes, a self-training strategy is employed. This involves creating a self-training set of unlabeled X-ray images from seen and unseen classes. The self-training is executed in two steps: initial inference and model fine-tuning. Initially, class probabilities for unlabeled images from the self-training set are computed using the trained MVSE network. Images are then selected for both seen and unseen classes based on the highest confidence scores. Subsequently, the model is fine-tuned with this selectively chosen data for each class. This refined model is then deployed for generalized zero-shot diagnosis of chest X-rays. During testing, for a given X-ray image, the model computes distances in both the X-ray and CT semantic spaces from the respective class signatures, dynamically balancing the importance of each branch to determine the final class probability. This model demonstrated robust generalization capabilities on the NIH Chest X-ray dataset (NIH), a hand-labeled subset of NIH dataset (NIH-900), Open-i dataset, PubMed Central dataset (PMC) and the CheXpert dataset.
@@
The Balanced Gradient Contribution (BGC) strategy is a training approach designed to manage the significant statistical differences between domains [152]. This method addresses the issue of large variance in gradients due to the distinct nature of data from each domain. In the context of DG, the BGC method could be employed to balance the learning from different domains by adjusting the contribution of gradients from each domain during the training process.
Elbatel et al. [90] integrated BGC into their Seamless Iterative Semi-supervised correction of imperfect labels (SISSI) [^27] https://github.com/marwankefah/SISSI, which trains object detection models with noisy and missing annotations. They introduced a range of image processing and deep learning methods to make iterative label correction. Using a domain adaptation strategy, they leveraged a source labeled dataset to enhance training on a target noisy dataset. Initially, they used a mixed-batch training with both training datasets to train a Faster R-CNN model using BGC [152], ensuring stable gradient directions. They used ADELE method to detect when the network starts memorizing the initial noisy annotations. Next, in the semi-supervised phase, they applied a label correction strategy using test-time augmentation and weighted box fusion techniques to produce confident bounding boxes.
@@
Robust loss function
Robust loss functions focus mainly on improving the loss to build robust DL network.
To address the problem of model overfitting due to label ambiguity and noisy labels, Sun et al. [91] proposed to use deep log-normal label distribution learning and focal loss. This approach is inspired by label distribution learning [153, 154], where an instance is assigned a label distribution, aiming to learn a mapping from instance to label distribution. For pneumoconiosis staging on chest radiographs, the authors modeled the label distribution using an asymmetric log-normal distribution.
@@
The KL Divergence loss was employed to enforce label distribution learning by measuring the distance between the label distribution ($y^{d}$) and the network prediction after the Softmax function. In addition, a regularization term, the cross-entropy loss, was added to strengthen the learning abilities of the model on unambiguous samples and handle subjective inconsistencies. The combination of KL divergence loss and cross-entropy loss forms the focal staging loss. To resolve optimization inconsistency when using these losses together, an instance-level drop parameter was introduced to skip samples with better predicted results during the optimization process.
Zhu et al. [80] employed the focal loss (Eq. 11) to improve training by emphasizing hard samples.
@@
where $q(y|x)$ is the predicted probability and $\gamma$ is a hyperparameter. $\gamma$ was set to 2 to reduce the relative loss for well-classified examples and focusing more on hard, misclassified one. This technique enhances the robustness of the model against noisy labels and ensures effective learning from challenging data.
Hu et al. [92] proposed a robust training method, Deep Supervised Network with a Self-Adaptive Auxiliary Loss (DSN-SAAL), for diagnosing imbalanced CT images. This framework integrates a novel loss function to address both the effects of data overlap between CT slices and noisy labels. To account for data overlap between CT slices, they adjusted the weight of samples in the Cross-Entropy (CE) loss function.
@@
where $q(y|x)$ is the classifier’s output and $p(y|x)$ is the ground-truth label distribution. $k_{i}$ is the number of samples in the $k^{th}$ class and $\alpha$ is a learnable parameter representing the effective sample factor to measure the ratio of the effective number of samples. To combat noisy labels, they introduced the Reverse Cross Entropy (RCE) loss:
@@
Here, $q(y|x)$ is used as the ground truth and $p(y|x)$ is the class probability of the outputs, hence the name reverse cross-entropy. Finally, the self-adaptive auxiliary loss combines the aforementioned losses (Equations 12 and 13) while adding a weighting hyperparameter $\beta$. This approach outperformed the state-of-the-art methods when evaluated on COVID19-Diag and three public COVID-19 diagnosis datasets.
To address annotation subjectivity, Yu et al. [93] proposed the Grading Cross Entropy (GCE) loss, designed to account for the feature continuity of disease grades and progression of disease grades. Misclassifications are more likely between adjacent grades than distant ones. The GCE loss is defined as follows:
@@
where $p(y_{i}|x)$ is the $i$-th element of the one-hot encoded label of the input $x$, $\mathbb{N}(i)$ the neighboring indexes of grade $i$, $q(y_{j}|x)$ denotes the $j$-th element of the model predictions and $w_{ij}$ represents the weight of grade $j$ with the annotated label $i$. This weighting system allows the GCE loss to be more flexible than the CE loss in handling noise by setting different weights to neighboring labels and the annotated label.
Hermoza et al. [76] tackled the problem of predicting survival time from medical images using both censored and uncensored data. They proposed an Early-Learning Regularization (ELR) loss, a regularization loss to manage noisy pseudo labels in survival time prediction. The ELR loss ensures continuous training for samples where the model’s prediction aligns with the temporal ensembling momentum (i.e., the “clean” pseudo-labeled samples) and ceases training for noisy pseudo-labeled samples. The ELR loss was expressed as follows:
@@
where $\tilde{z}_{i}^{c(e)}=\phi\tilde{z}_{i}^{c(e-1)}+(1-\phi)z_{i}^{c(e)}$ is the temporal ensembling momentum of the prediction ($z_{i}^{c}$) with $e$ denoting the training epoch and $\phi\in[0,1]$. $\sigma$ is the sigmoid function, $z^{c}$ is the model’s output, and $c$ denotes the number of classes.
Liu et al. [94] proposed a new training module called Non-Volatile Unbiased Memory (NVUM) [^28] https://github.com/FBLADL/NVUM, which stores a running average of model logits. They employed a regularization loss to minimize the differences between the model’s current logits and those from its initial learning phase. The model $f_{\theta}$ is trained on a noisy labeled dataset using binary cross-entropy loss, combined with the following regularization term:
@@
where $\tilde{z}^{c}$ stores an unbiased multi-label running average of the predicted logits of all training samples and employs the class prior distribution $\pi$ for updating. Initially, $\tilde{z}$ is initialized with zeros and is updated in every epoch as follows:
@@
where $\beta\in[0,1]$ is a hyperparameter controlling the volatility of the memory storage. $\beta$ was set to $0.9$ representing a non-volatile memory. This regularization enforces consistency between the current model logits and the logits produced at the beginning of the training, assuming robustness to noisy labels in early training. NVUM was evaluated on noisy multi-label imbalanced chest X-ray (CXR) training sets, formed by Chest-Xray14 and CheXpert, and tested on clean datasets OpenI and PadChest. The approach outperformed previous state-of-the-art classifiers with mean testing AUC of 0.8865 and 0.8555, respectively.
@@
Shi et al. [95] proposed a semi-supervised DL approach, Graph Temporal Ensembling (GTE), which leverages both labeled and unlabeled data while being robust to noisy labels. Inspired by Temporal Ensembling (TE) [155], GTE creates ensemble targets for feature and label predictions through Exponential Moving Average (EMA) to aggregate feature and label predictions from previous training epochs. Then, the ensemble targets within the same class are aggregated into clusters for further enhancement.The method also utilizes a consistency loss, which minimizes the discrepancy between the current predictions and the ensemble targets, to form consensus predictions under different configurations. The authors validated the proposed method with extensive experiments on lung and breast cancer datasets, achieving 90.5% and 89.5% image classification using 20% labeled patients on the two datasets, respectively.